# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 2 Some preliminaries
$Tor_{i}(\mathscr{D})=\{\textbf{w}\in\mathcal{T}_{m}^{n}:u^{i-1}(\textbf{w}+u\textbf{w}^{\prime})\in\mathscr{D}\text{ for some }\textbf{w}^{\prime}\in\mathscr{R}_{e,m}^{n}\}$
@@
Yadav and Sharma [31, 32, 33] counted all self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$ in the following three cases: (i) $p$ is an odd prime (i.e., the ring $\mathscr{R}_{e,m}$ is of odd characteristic), (ii) $p=2$ and $\mathfrak{s}=1$ (i.e., $\mathscr{R}_{e,m}$ is the quasi-Galois ring $\mathbb{F}_{2^{m}}[u]/\langle u^{e}\rangle$ of even characteristic), and (iii) $p=2$ and $\kappa=1$ (i.e., $\mathscr{R}_{e,m}$ is the Galois ring $GR(2^{e},m)$ of characteristic $2^{e}$ and cardinality $2^{em}$). When either $p$ is odd or $p=2$ and $\mathfrak{s}=1,$ they provided recursive methods for constructing self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{1},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ from a self-orthogonal code of length $n$ and dimension $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{\lceil\frac{e}{2}\rceil}$ over $\mathcal{T}_{m},$ where $\lambda_{1},\lambda_{2},\ldots,\lambda_{e}$ are non-negative integers satisfying $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-i+1}+\lambda_{e-i+2}+\cdots+\lambda_{i}\leq n$ for $\lceil\frac{e+1}{2}\rceil\leq i\leq e$ (see [32, Sec. 4] and [33, Sec. 5]). On the other hand, when $p=2$ and $\kappa=1,$ they provided a recursive method for constructing self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{1},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ from a doubly even code of length $n$ and dimension $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{\lceil\frac{e}{2}\rceil}$ over $\mathcal{T}_{m},$ where $\lambda_{1},\lambda_{2},\ldots,\lambda_{e}$ are non-negative integers satisfying $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-i+1}+\lambda_{e-i+2}+\cdots+\lambda_{i}\leq n$ for $\lceil\frac{e+1}{2}\rceil\leq i\leq e$ (see [31, Sec. 4]). It is worth noting that when $p=2$ and either $\mathfrak{s}=1$ or $\kappa=1,$ the recursive method provided in [33, Sec. 5] does not always lift a self-orthogonal code over $\mathcal{T}_{m}$ to a self-orthogonal code over $\mathscr{R}_{e,m}$ (see [31, Sec. 1] and [32, Sec. 3]). When $p=2$ and both $\mathfrak{s},\kappa\geq 2,$ the following example illustrates that the recursive methods provided in [31, Sec. 4] and [32, Sec. 4] do not always lift a self-orthogonal code over $\mathcal{T}_{m}$ and a doubly even code over $\mathcal{T}_{m}$ (see [31, Def. 2.1]) to a self-orthogonal code over $\mathscr{R}_{e,m},$ respectively.
@@
Let $\mathscr{R}_{8,2}=\frac{GR(8,2)[y]}{\langle y^{3}+2,4y^{2}\rangle}$ be a finite commutative chain ring with maximal ideal $\langle u\rangle$ of nilpotency index $8$ and $u^{3}+u^{6}=2$ in $\mathscr{R}_{8,2},$ where $u:=y+\langle y^{3}+2,4y^{2}\rangle.$ Let $n=4,$ $\lambda_{1}=1$ and $\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda_{5}=\lambda_{6}=\lambda_{7}=\lambda_{8}=0.$ Let $\mathcal{C}_{0}$ be a self-orthogonal code of length $4$ and dimension $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=1$ over $\mathcal{T}_{2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$. Note that $\pi_{1}(\textbf{d}\cdot\textbf{d})=\pi_{3}(\textbf{d}\cdot\textbf{d})=0$ for all $\textbf{d}\in\mathcal{C}_{0},$ where each $\textbf{d}\cdot\textbf{d}$ is viewed as an element of $\mathscr{R}_{8,2}.$
@@
Here, we will show that the code $\mathcal{C}_{0}$ can not be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ using the recursive construction method provided in Section 4 of Yadav and Sharma [32]. Towards this, let us consider a linear code $\mathcal{C}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&a_{1}&b_{1}&c_{1}\end{bmatrix}$, where $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}.$ Using the recursive construction method provided in Section 4 of [32], we see that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathcal{C}_{2}$ is a self-orthogonal code over $\mathscr{R}_{2,2}$ satisfying property $(\ast)$ (see [32, Sec. 4]), which holds if and only if $a_{1}^{2}+b_{1}^{2}+c_{1}^{2}=0.$ In particular, for $a_{1}=b_{1}=c_{1}=0,$ the code $\mathcal{C}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$ is a self-orthogonal code satisfying property $(\ast).$ Corresponding to this particular choice of the code $\mathcal{C}_{2}$, consider a linear code $\mathcal{C}_{4}$ of type $\{1,0,0,0\}$ and length $4$ over $\mathscr{R}_{4,2}$ with a generator matrix
@@
We will next show that the code $\mathcal{C}_{0}$ can not be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ using the recursive construction method provided in Section 4 of Yadav and Sharma [31]. For this, we first recall [31, Def. 2.1] that a self-orthogonal code $\mathscr{D}$ over $\mathcal{T}_{2}$ is said to be doubly even if it satisfies $\pi_{1}(\textbf{d}\cdot\textbf{d})=0$ for all $\textbf{d}\in\mathscr{D},$ where each $\textbf{d}\cdot\textbf{d}$ is viewed as an element of $\mathscr{R}_{8,2}$. One can easily see that $\pi_{1}(\textbf{v}\cdot\textbf{v})=0$ for all $\textbf{v}\in\mathcal{C}_{0},$ so the code $\mathcal{C}_{0}$ is doubly even. Now, let us consider a linear code $\mathcal{D}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&f_{1}&g_{1}&h_{1}\end{bmatrix}$, where $f_{1},g_{1},h_{1}\in\mathcal{T}_{2}.$ By the recursive construction method provided in Section 4 of [31], we see that the code $\mathcal{D}_{2}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathcal{D}_{2}$ is a $\lambda_{1}$-doubly even self-orthogonal code over $\mathscr{R}_{2,2}$ (see [31, Def. 2.1]), which holds if and only if $f_{1}^{2}+g_{1}^{2}+h_{1}^{2}\equiv 0\pmod{u}$. In particular, for $f_{1}=g_{1}=h_{1}=0,$ we see that the code $\mathcal{D}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$ is a $\lambda_{1}$-doubly even self-orthogonal code over $\mathscr{R}_{2,2}.$ Corresponding to this particular choice of the code $\mathcal{D}_{2}$, let us consider a linear code $\mathcal{D}_{4}$ of type $\{1,0,0,0\}$ and length $4$ over $\mathscr{R}_{4,2}$ with a generator matrix
@@
From now on, we assume, throughout this paper, that $p=2$ and the integer $\kappa\geq 3$ is odd, i.e., the element $2\in\langle u^{\kappa}\rangle\setminus\langle u^{\kappa+1}\rangle$ with $\kappa\geq 3$ being odd. Accordingly, we assume, from now on, that $2\equiv u^{\kappa}(\eta_{0}+u\eta_{1}+u^{2}\eta_{2}+\cdots+u^{e-1-\kappa}\eta_{e-1-\kappa})\pmod{u^{e}}$, where $\eta_{0}\in\mathcal{T}_{m}\setminus\{0\}$ and $\eta_{1},\eta_{2},\ldots,\eta_{e-1-\kappa}\in\mathcal{T}_{m}.$ In the following section, we will provide a recursive method for lifting a self-orthogonal code over $\mathcal{T}_{m}$ to a self-orthogonal code over $\mathscr{R}_{e,m}$ when $\kappa\geq 3$ is odd. We will consider the case when $\kappa$ is even in a subsequent work [30]. To do this, let us define $\mathrm{f}_{w}=\big\lfloor\frac{w}{2}\big\rfloor$ and $\mathrm{c}_{w}=\big\lceil\frac{w}{2}\big\rceil$ for any positive integer $w.$ Let us define $\theta_{e}=0$ if $e$ is even, while $\theta_{e}=1$ if $e$ is odd, and let us define the numbers $s=\mathrm{f}_{e}=\big\lfloor\frac{e}{2}\big\rfloor$ and $\kappa_{1}=\mathrm{f}_{\kappa}=\big\lfloor\frac{\kappa}{2}\big\rfloor=\frac{\kappa-1}{2}$ for our convenience. Unless otherwise stated, we will adhere to the notations introduced in Section 2.
@@
By carrying out direct computations, one can easily observe that linear codes over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ coincide with linear codes over $\mathscr{R}_{\ell,m}$ satisfying property $(\ast)$ when $\mathfrak{s}=1$ [32, Sec. 4]. Next, for positive integers $v$ and $w\leq e,$ let $(\mathtt{H})_{v,w}$ denote the block matrix whose $(i,j)$-th block is the matrix $\mathtt{H}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq v$ and $w\leq j\leq e.$ Further, if $\mathscr{D}$ is a self-orthogonal code of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$, then we see, by Lemma 2.2, that the torsion code $Tor_{s+\theta_{e}}(\mathscr{D})$ is also a self-orthogonal code of length $n$ and dimension $\Lambda_{s+\theta_{e}}$ over $\mathcal{T}_{m}(=\overline{\mathscr{R}}_{e,m}).$ We now make the following observation.
@@
$G=\begin{bmatrix}\mathtt{I}_{\lambda_{1}}&\mathtt{B}_{1,1}&\mathtt{B}_{1,2}&\cdots&\mathtt{B}_{1,s+\theta_{e}-1}&\mathtt{B}_{1,s+\theta_{e}}&\cdots&\mathtt{B}_{1,e}\\ \mathbf{0}&\mathtt{I}_{\lambda_{2}}&\mathtt{B}_{2,2}&\cdots&\mathtt{B}_{2,s+\theta_{e}-1}&\mathtt{B}_{2,s+\theta_{e}}&\cdots&\mathtt{B}_{2,e}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\cdots&\mathtt{I}_{\lambda_{s+\theta_{e}}}&\mathtt{B}_{s+\theta_{e},s+\theta_{e}}&\cdots&\mathtt{B}_{s+\theta_{e},e}\end{bmatrix}$
whose columns are partitioned into blocks of sizes $\lambda_{1},\lambda_{2},\ldots,\lambda_{e+1},$ the matrix $\mathtt{I}_{\lambda_{i}}$ is the $\lambda_{i}\times\lambda_{i}$ identity matrix over $\mathcal{T}_{m},$ $\mathtt{B}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq s+\theta_{e}$ and $i\leq j\leq e,$ and the matrices $(\mathtt{B})_{s,s+\theta_{e}+1}$, $(\mathtt{B})_{s-1,s+\theta_{e}+2},\ldots,(\mathtt{B})_{2,e-1},$ $\mathtt{B}_{1,e}$ are of full row ranks.
@@
$\mathcal{G}_{e}=\begin{bmatrix}\mathtt{W}^{(e)}_{1}\\ u\mathtt{W}^{(e)}_{2}\\ u^{2}\mathtt{W}^{(e)}_{3}\\ \vdots\\ u^{e-1}\mathtt{W}^{(e)}_{e}\\ \end{bmatrix}$, (3.1)
@@
      Now, for $i\in\{2,4,\ldots,\ell-\theta_{e}\},$ as $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ we see, by Lemma 2.1, that $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, from which one can easily observe that Diag([W^(ℓ)]_γ_ℓ+1-f_i[W^(ℓ)]_γ_ℓ+1-f_i^t)Diag([W^(ℓ+i)]_γ_ℓ+1-f_i[W^(ℓ+i)]_γ_ℓ+1-f_i^t)0(modu^ℓ+i). Further, for $j\in\{\ell-1,\ell+1,\ldots,\kappa-2+\theta_{e}\},$ since $\mathscr{D}_{\ell+j+1+\theta_{e}}$ is a self-orthogonal code over $\mathscr{R}_{\ell+j+1+\theta_{e},m},$ by Lemma 2.1 again, we get
@@
      We further observe, for $1\leq a\leq\gamma_{\ell}+1-\mathtt{f}_{j+2},$ that the $(a,a)$-th entry, say $L_{a}$, of the matrix $\mathtt{Diag}\big([\mathtt{W}^{(\ell+j+1+\theta_{e})}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}[\mathtt{W}^{(\ell+j+1+\theta_{e})}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}^{t}\big)$ satisfies

### v2

@@ 2.4 Torsion codes over $\mathcal{T}_{m}$ and self-orthogonal and self-dual codes over $\mathscr{R}_{e,m}$
$Tor_{i}(\mathscr{D})=\{\textbf{w}\in\mathcal{T}_{m}^{n}:u^{i-1}(\textbf{w}+u\textbf{w}^{\prime})\in\mathscr{D}\text{ for some }\textbf{w}^{\prime}\in\mathscr{R}_{e,m}^{n}\}\vskip-5.69054pt$
@@
2.5 Limitations of existing constructions
Yadav and Sharma [34, 35, 36] counted all self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$ in the following three cases: (i) $p$ is an odd prime (i.e., the ring $\mathscr{R}_{e,m}$ is of odd characteristic), (ii) $p=2$ and $\mathfrak{s}=1$ (i.e., $\mathscr{R}_{e,m}$ is the quasi-Galois ring $\mathbb{F}_{2^{m}}[u]/\langle u^{e}\rangle$ of even characteristic), and (iii) $p=2$ and $\kappa=1$ (i.e., $\mathscr{R}_{e,m}$ is the Galois ring $GR(2^{e},m)$ of characteristic $2^{e}$ and cardinality $2^{em}$). In view of Remark 2.1, let $\lambda_{1},\lambda_{2},\ldots,\lambda_{e}$ be non-negative integers satisfying $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-i+1}+\lambda_{e-i+2}+\cdots+\lambda_{i}\leq n$ for $\lceil\frac{e+1}{2}\rceil\leq i\leq e.$ When either $p=2$ and $\mathfrak{s}=1$ or $p$ is odd, Yadav and Sharma [35, Sec. 4] and [36, Sec. 5] provided recursive methods for constructing self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{1},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ from a self-orthogonal code of length $n$ and dimension $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{\lceil\frac{e}{2}\rceil}$ over $\mathcal{T}_{m}.$ On the other hand, when $p=2$ and $\kappa=1,$ Yadav and Sharma [34, Sec. 4] provided a recursive method for constructing self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{1},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$ from a doubly even code of length $n$ and dimension $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{\lceil\frac{e}{2}\rceil}$ over $\mathcal{T}_{m}.$ It is worth noting that when $p=2$ and either $\mathfrak{s}=1$ or $\kappa=1,$ the recursive method provided in [36, Sec. 5] does not always lift a self-orthogonal code over $\mathcal{T}_{m}$ to a self-orthogonal code over $\mathscr{R}_{e,m}$ (see [34, Sec. 1] and [35, Sec. 3]). When $p=2$ and both $\mathfrak{s},\kappa\geq 2,$ the following example illustrates that the recursive methods provided in [34, Sec. 4] and [35, Sec. 4] do not always lift a self-orthogonal code over $\mathcal{T}_{m}$ and a doubly even code over $\mathcal{T}_{m}$ (see [34, Def. 2.1]) to a self-orthogonal code over $\mathscr{R}_{e,m},$ respectively.
@@ 2.5 Limitations of existing constructions
Let $\mathscr{R}_{8,2}=\frac{GR(8,2)[y]}{\langle y^{3}+2,4y^{2}\rangle}$ be a finite commutative chain ring with maximal ideal $\langle u\rangle$ of nilpotency index $8,$ where $u:=y+\langle y^{3}+2,4y^{2}\rangle\in\mathscr{R}_{8,2}.$ Note that $u^{3}+u^{6}=2$ in $\mathscr{R}_{8,2}.$ Let $n=4,$ $\lambda_{1}=1$ and $\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda_{5}=\lambda_{6}=\lambda_{7}=\lambda_{8}=0.$ Let $\mathcal{C}_{0}$ be a self-orthogonal code of length $4$ and dimension $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=1$ over $\mathcal{T}_{2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$. Note that $\pi_{1}(\textbf{d}\cdot\textbf{d})=\pi_{3}(\textbf{d}\cdot\textbf{d})=0$ for all $\textbf{d}\in\mathcal{C}_{0},$ where each $\textbf{d}\cdot\textbf{d}$ is viewed as an element of $\mathscr{R}_{8,2}.$
@@
Here, we will show that the code $\mathcal{C}_{0}$ can not be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ using the recursive construction method provided in Section 4 of Yadav and Sharma [35]. Towards this, let us consider a linear code $\mathcal{C}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&a_{1}&b_{1}&c_{1}\end{bmatrix}$, where $a_{1},b_{1},c_{1}\in\mathcal{T}_{2}.$ Using the recursive construction method provided in Section 4 of [35], we see that the code $\mathcal{C}_{0}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathcal{C}_{2}$ is a self-orthogonal code over $\mathscr{R}_{2,2}$ satisfying property $(\ast)$ (see [35, Sec. 4]), which holds if and only if $a_{1}^{2}+b_{1}^{2}+c_{1}^{2}=0.$ In particular, for $a_{1}=b_{1}=c_{1}=0,$ the code $\mathcal{C}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$ is a self-orthogonal code satisfying property $(\ast).$ Corresponding to this particular choice of the code $\mathcal{C}_{2}$, consider a linear code $\mathcal{C}_{4}$ of type $\{1,0,0,0\}$ and length $4$ over $\mathscr{R}_{4,2}$ with a generator matrix
@@
We will next show that the code $\mathcal{C}_{0}$ can not be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ and length $4$ over $\mathscr{R}_{8,2}$ using the recursive construction method provided in Section 4 of Yadav and Sharma [34]. For this, we first recall [34, Def. 2.1] that a self-orthogonal code $\mathscr{D}$ over $\mathcal{T}_{2}$ is said to be doubly even if it satisfies $\pi_{1}(\textbf{d}\cdot\textbf{d})=0$ for all $\textbf{d}\in\mathscr{D},$ where each $\textbf{d}\cdot\textbf{d}$ is viewed as an element of $\mathscr{R}_{8,2}$. One can easily see that $\pi_{1}(\textbf{v}\cdot\textbf{v})=0$ for all $\textbf{v}\in\mathcal{C}_{0},$ so the code $\mathcal{C}_{0}$ is doubly even. Now, let us consider a linear code $\mathcal{D}_{2}$ of type $\{1,0\}$ and length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}+u\begin{bmatrix}0&f_{1}&g_{1}&h_{1}\end{bmatrix}$, where $f_{1},g_{1},h_{1}\in\mathcal{T}_{2}.$ By the recursive construction method provided in Section 4 of [34], we see that the code $\mathcal{D}_{2}$ can be lifted to a self-orthogonal code of type $\{1,0,0,0,0,0,0,0\}$ over $\mathscr{R}_{8,2}$ if and only if $\mathcal{D}_{2}$ is a $\lambda_{1}$-doubly even self-orthogonal code over $\mathscr{R}_{2,2}$ (see [34, Def. 2.1]), which holds if and only if $f_{1}^{2}+g_{1}^{2}+h_{1}^{2}\equiv 0\pmod{u}$. In particular, for $f_{1}=g_{1}=h_{1}=0,$ we see that the code $\mathcal{D}_{2}$ of length $4$ over $\mathscr{R}_{2,2}$ with a generator matrix $\begin{bmatrix}1&1&1&1\end{bmatrix}$ is a $\lambda_{1}$-doubly even self-orthogonal code over $\mathscr{R}_{2,2}.$ Corresponding to this particular choice of the code $\mathcal{D}_{2}$, let us consider a linear code $\mathcal{D}_{4}$ of type $\{1,0,0,0\}$ and length $4$ over $\mathscr{R}_{4,2}$ with a generator matrix
@@
2.6 Notations and preliminary setup
From now on, we assume, throughout this paper, that $p=2$ and the integer $\kappa\geq 3$ is odd, i.e., the element $2\in\langle u^{\kappa}\rangle\setminus\langle u^{\kappa+1}\rangle$ with $\kappa\geq 3$ being odd. Accordingly, we assume, from now on, that $2\equiv u^{\kappa}(\eta_{0}+u\eta_{1}+u^{2}\eta_{2}+\cdots+u^{e-1-\kappa}\eta_{e-1-\kappa})\pmod{u^{e}}$, where $\eta_{0}\in\mathcal{T}_{m}\setminus\{0\}$ and $\eta_{1},\eta_{2},\ldots,\eta_{e-1-\kappa}\in\mathcal{T}_{m}.$ In the following section, we will provide a recursive method for lifting a self-orthogonal code over $\mathcal{T}_{m}$ to a self-orthogonal code over $\mathscr{R}_{e,m}$ when $\kappa\geq 3$ is odd; the case when $\kappa$ is even will be addressed in a subsequent work [33]. For notational convenience, we define $\mathrm{f}_{w}=\big\lfloor\frac{w}{2}\big\rfloor$ and $\mathrm{c}_{w}=\big\lceil\frac{w}{2}\big\rceil$ for any positive integer $w,$ $\theta_{e}=0$ if $e$ is even and $\theta_{e}=1$ if $e$ is odd, $s=\mathrm{f}_{e}=\big\lfloor\frac{e}{2}\big\rfloor,$ and $\kappa_{1}=\mathrm{f}_{\kappa}=\big\lfloor\frac{\kappa}{2}\big\rfloor=\frac{\kappa-1}{2}.$ Unless otherwise stated, we will adhere to the notations introduced in Section 2.
@@ 2.6 Notations and preliminary setup
By carrying out direct computations, one can easily observe that linear codes over $\mathscr{R}_{\ell,m}$ satisfying property $(\mathfrak{P})$ coincide with linear codes over $\mathscr{R}_{\ell,m}$ satisfying property $(\ast)$ when $\mathfrak{s}=1$ [35, Sec. 4]. Next, for positive integers $v$ and $w\leq e,$ let $(\mathtt{H})_{v,w}$ denote the block matrix whose $(i,j)$-th block is the matrix $\mathtt{H}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq v$ and $w\leq j\leq e.$ Further, if $\mathscr{D}$ is a self-orthogonal code of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m}$, then we see, by Lemma 2.2, that the torsion code $Tor_{s+\theta_{e}}(\mathscr{D})$ is also a self-orthogonal code of length $n$ and dimension $\Lambda_{s+\theta_{e}}$ over $\mathcal{T}_{m}(=\overline{\mathscr{R}}_{e,m}).$ We now make the following observation.
@@
[…]
where the columns of $G$ are partitioned into blocks of sizes $\lambda_{1},\lambda_{2},\ldots,\lambda_{e+1},$ $\mathtt{I}_{\lambda_{i}}$ denotes the $\lambda_{i}\times\lambda_{i}$ identity matrix over $\mathcal{T}_{m},$ $\mathtt{B}_{i,j}\in\mathcal{M}_{\lambda_{i}\times\lambda_{j+1}}(\mathcal{T}_{m})$ for $1\leq i\leq s+\theta_{e}$ and $i\leq j\leq e,$ and the matrices $(\mathtt{B})_{s,s+\theta_{e}+1}$, $(\mathtt{B})_{s-1,s+\theta_{e}+2},\ldots,(\mathtt{B})_{2,e-1},$ $\mathtt{B}_{1,e}$ are of full row ranks.
@@
[…]
@@
Now, for $i\in\{2,4,\ldots,\ell-\theta_{e}\},$ as $\mathscr{D}_{\ell+i}$ is a self-orthogonal code over $\mathscr{R}_{\ell+i,m},$ we see, by Lemma 2.1, that $\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$, from which one can easily observe that
$\mathtt{Diag}\big([\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathtt{Diag}\big([\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}[\mathtt{W}^{(\ell+i)}]_{\gamma_{\ell}+1-\mathrm{f}_{i}}^{t}\big)\equiv\mathbf{0}\pmod{u^{\ell+i}}$.
Further, for $j\in\{\ell-1,\ell+1,\ldots,\kappa-2+\theta_{e}\},$ since $\mathscr{D}_{\ell+j+1+\theta_{e}}$ is a self-orthogonal code over $\mathscr{R}_{\ell+j+1+\theta_{e},m},$ by Lemma 2.1 again, we get
@@
We further observe, for $1\leq a\leq\gamma_{\ell}+1-\mathrm{f}_{j+2},$ that the $(a,a)$-th entry, say $L_{a}$, of the matrix $\mathtt{Diag}\big([\mathtt{W}^{(\ell+j+1+\theta_{e})}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}[\mathtt{W}^{(\ell+j+1+\theta_{e})}]_{\gamma_{\ell}+1-\mathrm{f}_{j+2}}^{t}\big)$ satisfies